OpenStudy (anonymous):

how do you differentiate the square root of 2x-3?

OpenStudy (jhannybean):

chain rule.

OpenStudy (anonymous):

\[\frac{d}{dx}[\sqrt{f(x)}]=\frac{f'(x)}{2\sqrt{f(x)}}\]

OpenStudy (solomonzelman):

At first, do you know the derivative of a sqrt(x) ?

OpenStudy (solomonzelman):

\(\large\color{black}{ \sqrt{x} }\) is same as \(\large\color{black}{ x^{1/2} }\). So apply the power rule, \(\large\color{black}{ \frac{d}{dx}~x^n ~=n~\times x^{n-1} }\)

OpenStudy (solomonzelman):

I'll give you examples if you want to.

OpenStudy (solomonzelman):

I keep getting disconnected. god dang it.

OpenStudy (solomonzelman):

1) Find the derivative of \(\large\color{black}{ \sqrt{x^2-4x+5} }\). First derive the function as if it was just a square root of \(\large\color{black}{ \sqrt{x} }\) , then using a chain rule, mutliply times the derivative of the inner function, of \(\large\color{black}{ x^2-4x+5 }\). the derivative of \(\large\color{black}{ \sqrt{x} }\) is \(\large\color{black}{ \frac{d}{dx}~\sqrt{x}=\frac{d}{dx}x^{1/2}=\frac{1}{2}\times x^{1/2~~~-1}=\frac{1}{2}\times x^{-1/2}=\frac{1}{2x^{1/2}}=\frac{1}{2\sqrt{x}}}\) the derivative of \(\large\color{black}{ x^2-4x+5 }\) (applying power rule to each term) is: \(\large\color{black}{ 2x-4 }\). So, we get this: \(\Large\color{black}{ \frac{d}{dx}(\sqrt{x^2-4x+5}~~)=\frac{1}{2\sqrt{x^2-4x+5}}\times (2x-4) }\) this simplifies to, \(\Large\color{blue}{ \frac{d}{dx}(\sqrt{x^2-4x+5}~~)=\frac{(2x-4)}{2\sqrt{x^2-4x+5}} }\) and to completely simplify: \(\Large\color{blue}{ \frac{d}{dx}(\sqrt{x^2-4x+5}~~)=2\frac{(x-2)}{2\sqrt{x^2-4x+5}} }\) \(\Large\color{red}{ \frac{d}{dx}(\sqrt{x^2-4x+5}~~)=\frac{(x-2)}{\sqrt{x^2-4x+5}} }\) and this is the final answer.